# Generalization under Byzantine & Poisoning Attacks: Tight Stability Bounds in Robust Distributed Learning

### v1

@@ D.1 Expectation bound for the honest gradients’ empirical covariance trace and spectral norm
$\mathbb{E}_{\mathcal{A}}[\|G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{SGD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}]\leq\eta_{G^{\mathrm{SGD}}_{\gamma_{t}}}\mathbb{E}_{\mathcal{A}}[\|\theta_{t}-\theta^{\prime}_{t}\|_{2}]+\frac{2\gamma_{t}C}{(n-f)m}$ (10)
Let define $\sigma^{F}_{\mathrm{SGD}}=2C\left(\frac{1}{(n-f)m}+\sqrt{\kappa}\right)$, we then resolve the recursion by incorporating the impact of the Byzantine failure attack as demonstrated in the previous proof sketch.
@@
    $\mathbb{E}_{\mathcal{A}}[\|\theta_{t+1}-\theta_{t+1}^{\prime}\|_{2}]\leq\mathbb{E}_{\mathcal{A}}[\|\theta_{t}-\theta_{t}^{\prime}\|]+\gamma_{t}\sigma^{F}_{\mathrm{SGD}}$
@@
    $\mathbb{E}_{\mathcal{A}}\|\theta_{t+1}-\theta_{t+1}^{\prime}\|_{2}\leq(1-\gamma\mu)\mathbb{E}_{\mathcal{A}}\|\theta_{t}-\theta_{t}^{\prime}\|+\gamma\sigma^{F}_{\mathrm{SGD}}$
    $\mathbb{E}_{\mathcal{A}}\|\theta_{T}-\theta_{T}^{\prime}\|_{2}$ $\leq\gamma\sigma^{F}_{\mathrm{SGD}}\sum_{t=0}^{T-1}{\left(1-\gamma\mu\right)}^{t}=\frac{\sigma^{F}_{\mathrm{SGD}}}{\mu}\left[1-{\left(1-\gamma\mu\right)}^{T}\right]\leq\frac{\sigma^{F}_{\mathrm{SGD}}}{\mu}$
  • (iii)
    If the loss function is Lipschitz, smooth, and nonconvex, then considering $t_{0}\in\{0,\ldots,T-1\}$, for all $t\in\{t_{0},\ldots,T-1\}$:
    $\mathbb{E}_{\mathcal{A}}[\delta_{t+1}|\delta_{t_{0}}=0]$ $\leq(1+\gamma_{t}L)\mathbb{E}_{\mathcal{A}}[\delta_{t}|\delta_{t_{0}}=0]+\gamma_{t}\sigma^{F}_{\mathrm{SGD}}$
    $\underset{1+x\leq e^{x}}{\leq}e^{\gamma_{t}L}\mathbb{E}_{\mathcal{A}}[\delta_{t}|\delta_{t_{0}}=0]+\gamma_{t}\sigma^{F}_{\mathrm{SGD}}$
    Let $c>0$ and consider a monotonically non-increasing learning rate $\gamma_{t}\leq\frac{c}{Lt}$, We proceed as follows:
    $\mathbb{E}_{\mathcal{A}}[\delta_{T}|\delta_{t_{0}}=0]\leq\frac{c\sigma^{F}_{\mathrm{SGD}}}{L}\sum_{t=t_{0}}^{T-1}\frac{1}{t}\prod_{s=t+1}^{T-1}e^{\frac{c}{s}},$
    where:
    $\sum_{t=t_{0}}^{T-1}\frac{1}{t}\prod_{s=t+1}^{T-1}e^{\frac{c}{s}}=\sum_{t=t_{0}}^{T-1}\frac{1}{t}e^{c\sum_{s=t+1}^{T-1}\frac{1}{s}}\leq\sum_{t=t_{0}}^{T-1}\frac{1}{t}e^{c\log(\frac{T}{t})}=T^{c}\sum_{t=t_{0}}^{T-1}t^{-c-1}\leq\frac{1}{c}{\left(\frac{T}{t_{0}}\right)}^{c}.$

### v2

@@ C.1 Expectation bounds for honest gradients’ empirical spectral norm
$\mathbb{E}_{\mathcal{A}}[\|G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{SGD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}]\leq\eta_{G^{\mathrm{SGD}}_{\gamma_{t}}}\mathbb{E}_{\mathcal{A}}[\|\theta_{t}-\theta^{\prime}_{t}\|_{2}]+\frac{2\gamma_{t}C}{(n-f)m}$
Let define $\sigma^{F}_{\mathrm{SGD}}=2C\left(\frac{1}{(n-f)m}+\sqrt{\kappa}\right)$, we then resolve the recursion, like in the $\mathrm{GD}$ case above, by incorporating the impact of the Byzantine failures as demonstrated in the previous proof sketch:
$\mathbb{E}_{\mathcal{A}}[\|\theta_{t+1}-\theta^{\prime}_{t+1}\|_{2}]\leq\eta_{G^{\mathrm{SGD}}_{\gamma_{t}}}\mathbb{E}_{\mathcal{A}}[\|\theta_{t}-\theta^{\prime}_{t}\|_{2}]+\gamma_{t}\sigma^{F}_{\mathrm{SGD}}.$ (14)
This conclude the proof. ∎
C.3 Proof of Theorem 3.2
Theorem 3.2.
Consider the setting described in Section 2 under Byzantine failures. Let $\mathcal{A}=\mathrm{GD}$, with $\mathrm{SMEA}$. Suppose $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations with the learning rate $\gamma$. Then for both regime stated below, there exist $\ell\in{\mathbb{R}}^{\Theta\times\mathcal{Z}}$ such that $\forall z\in\mathcal{Z},\ell(\cdot;z)$ is $C$-Lipschitz, $L$-smooth and convex, and a pair of neighboring datasets such that the uniform stability is lower bounded by
@@
    in the regime of moderate to high fractions of misbehaving workers $\frac{n}{3}\leq f<\frac{n}{2}$,
    $\Omega\left(\gamma C^{2}T\left(\sqrt{\frac{f}{n-2f}}+1+\frac{1}{(n-f)m}\right)\right).$
@@
    in the regime of low to moderate fractions of misbehaving workers $f<\frac{n}{3}$,
    $\Omega\left(\gamma C^{2}T\frac{f}{n-2f}\right).$
Proof.
The lower bound construction relies on two key elements. First, we maximize the variance among the honest gradients by dividing them into two equal subgroups positioned at the boundary of the Lipschitz constraint. Second, in the regime with a high proportion of Byzantine workers, this setup enables the adversary to completely discard one of the honest subgroups and even outweigh the influence of the other. The extent to which the adversary can bias the aggregate away from the true direction (which is zero by construction) depends on the robustness of the aggregation rule.
In what follows, we provide the proof for the case $d=1$. The result extends naturally to higher dimensions by embedding the one-dimensional construction into $\mathbb{R}^{d}$. We arbitrarily refer to one trajectory as unperturbed, with parameter $\theta_{t}$, and the other as perturbed, with $\theta^{\prime}_{t}$. Let $C,L\in\mathbb{R}^{*}_{+}$, we define the following loss function:
$\text{For }\theta\in\mathbb{R},\text{ }z\in[-C,C],\quad\ell(\theta;z)=z\theta,\quad\ell^{\prime}(\theta;z)=z,\quad\ell^{\prime\prime}(\theta;z)=0.$
By construction, the loss function is convex, $C$-Lipschitz and $L$-smooth. Let $m,n,f\in\mathbb{N}^{*}$ such that $\frac{n}{3}\leq f<\frac{n}{2}$. Let $\gamma$ the learning rate and assume $\theta_{0}=\theta_{0}^{\prime}=0$.
(i) If nfn-f is even and $f\geq\frac{n}{3}$.
We define the following pair of neighboring datasets:
$\begin{split}S=\{D_{1},\ldots,D_{n-f}\},&\quad S^{\prime}=S\setminus\{D_{1}\}\cup\{D_{1}\setminus\{z^{(1,1)}\}\cup\{z^{\prime(1,1)}\}\}\\ \text{ with }\forall i\in\{1,\ldots,n-f\},&\quad D_{i}=\{z^{(i,1)},\ldots,z^{(i,m)}\}\\ \text{ and }\forall j\in\{1,\ldots,m\},&\quad i\in\mathcal{H}_{+}=\{1,\ldots,\frac{n-f}{2}\},|\mathcal{H}_{+}|=\frac{n-f}{2}\!\!:z^{(i,j)}=C\text{ and }z^{\prime(1,1)}=0.\\ &\quad i\in\mathcal{H}_{-}=\{\frac{n-f}{2}+1,\ldots,n-f\},|\mathcal{H}_{-}|=\frac{n-f}{2}\!\!:z^{(i,j)}=-C.\end{split}$
For $t\in\{0,\ldots,T-1\}$, for $i\in\mathcal{H}_{-}$, $g^{(i)}_{t}=g^{(i)\prime}_{t}=-C$, for $i\in\mathcal{H}_{+}\setminus\{1\}$, $g^{(i)}_{t}=g^{(i)\prime}_{t}=C$, $g_{t}^{(1)}=C$ and $g_{t}^{(1)\prime}=\frac{m-1}{m}C<C$. We consider $f$ Byzantine workers, and let denote $\mathcal{B}=\{n-f+1,\ldots,n\}$. Since they can observe all communications, they may adapt their strategy based on the message sent by worker $1$. We define their behavior as follows:
if $g_{1}\leq C$ then
 return an arbitrarily high value (for instance $n\beta$)
else
 return $\beta$ (to be specified below)
Algorithm 1 Byzantine worker behavior
In the first case, the subset selected by the $\mathrm{SMEA}$ aggregation rule consists entirely of honest workers. Consequently, we have $\theta_{T}^{\prime}=\frac{\gamma CT}{m(n-f)}$. In the second case, we choose
$\beta=\left(1+\frac{1}{2}\frac{n-f}{\sqrt{f(n-2f)}}\right)C,$
such that the $\mathrm{SMEA}$ aggregation rule selects the subset $\text{Select}_{n-2f}(\mathcal{H}_{+})\cup\mathcal{B}$, where $\text{Select}_{k}(E)$ denote any subset of size $k$ from a set $E$, i.e., all the Byzantine workers together with n2fn-2f workers from $\mathcal{H}_{+}$. This is made possible because: